OpenStudy (anonymous):

A school fair ticket costs $8 per adult and $1 per child. On a certain day, the total number of adults (a) and children (c) who went to the fair was 30, and the total money collected was $100. Which of the following options represents the number of children and the number of adults who attended the fair that day, and the pair of equations that can be solved to find the numbers?

OpenStudy (anonymous):

A. 20 children and 10 adults Equation 1: a + c = 30 Equation 2: 8a + c = 100 B. 10 children and 20 adults Equation 1: a + c = 30 Equation 2: 8a – c = 100 C. 10 children and 20 adults Equation 1: a + c = 30 Equation 2: 8a + c = 100 D. 20 children and 10 adults Equation 1: a + c = 30 Equation 2: 8a – c = 100

OpenStudy (anonymous):

plz help me any one!!!

OpenStudy (ybarrap):

So the amount of money you make for adults depends on the number of adults, a, and the cost of their ticket: 8*a Similarly for children 1*c The total amount made for both adults and children then is 8a + 1c The total amount of people if the number of children plus the number of adults a + c Make sense?

OpenStudy (anonymous):

so is it A?

OpenStudy (ybarrap):

You are told that the total number of people on a certain day was 30 and they mad $100. Use the information I just gave you and equate dollars with the components that make up dollars and number, with the components that make of total number of people.

OpenStudy (ybarrap):

If you think the equations are Equation 1: a + c = 30 Equation 2: 8a + c = 100 Then, solve for a and c to see whether it's A or C

OpenStudy (anonymous):

K

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

C. { a + c = 30, 8 a + c = 100 } // Solve { a = 10, c = 20 }